# BEACON: Budget-Aware Entity Matching Across Domains

### 12 Mar 2026 v1

@@ Abstract
Entity Matching (EM)–the task of determining whether two data records refer to the same real-world entity–is a core task in data integration. Recent advances in deep learning have set a new standard for EM, particularly through fine-tuning Pretrained Language Models (PLMs) and, more recently, Large Language Models (LLMs). However, fine-tuning typically requires large amounts of labeled data, which are expensive and time-consuming to obtain. In the context of e-commerce matching, labeling scarcity varies widely across domains, raising the question of how to intelligently train accurate domain-specific EM models with limited labeled data. In this work we assume users have only limited amount of labels for a specific target domain but have access to labeled data from other domains. We introduce BEACON, a distribution-aware, budget-aware framework for low-resource EM across domains. BEACON leverages the insight that embedding representations of pairwise candidate matches can guide the effective selection of out-of-domain samples under limited in-domain supervision. We conduct extensive experiments across multiple domain-partitioned datasets derived from established EM benchmarks, demonstrating that BEACON consistently outperforms state-of-the-art methods under different training budgets.
@@ 1 Introduction
Figure 1 shows three entity pairs produced by blocking, with each pair comparing two product records from e-commerce datasets. The samples span multiple product categories and illustrate the types of candidate pairs that serve as inputs to EM models. A solid green line between records indicates a match, while a red dotted line indicates a non-match. Samples (a) and (c) represent matches, where both entities in the pair refer to the same product (though this is not known during matching). For example, sample (a) compares the same RAM stick, and sample (c) compares the same power supply. Conversely, sample (b) is a non-match: although both products are baseball caps from the same brand, they are distinct items. The record data is incomplete–as can be seen with missing ”brand” values–highlighting the challenges of real-world data systems and EM in practice.
Figure 2: Entity matching samples from Figure 1 grouped by the product category domain. The goal is to select a pairwise sample to complement the selected training set for EM in the ”Computers” product category.
@@
Suppose we aim to train a model for the “Computers” category. Figure 2 shows the same three pairwise samples from Figure 1, now labeled by product category. Note that any pairs spanning different product categories are automatically removed during blocking, allowing the remaining data to be partitioned into distinct category-specific groups. To build a training set for “Computers,” we first include all in-domain pairs, as they are the most likely be relevant to our model. In our toy example, only pair (a) belongs to the domain and is selected first. If room remains in the budget after including all in-domain samples, we need to decide which out-of-domain examples are most “beneficial”.
@@
  • Experimental Evaluation. We adapt EM PLMs (e.g., DITTO [27]), recent EM domain adaptation frameworks (e.g., MFSN [51]), recent low-resource EM methods (e.g., Battleship [19] and PromptEM [56]), and recent LLM approaches (e.g., Jellyfish [61, 62]) to establish strong baselines for EMAD, over which we conduct extensive experiments showing that BEACON outperforms approaches across different budget constraints. For this evaluation, we derive nine domain-partitioned datasets from WDC [38], as well as two additional datasets derived from the WDC Product Corpus [41] and Abt-Buy [25], which we also release to support future research on EMAD. We further provide Open-source Code and Data. [^2] https://github.com/nbpulsone/BEACON
@@ 2 Related Work
While our work shares the high-level goal of improving EM under domain shifts, there are two important distinctions that highlight the novelty of our problem setting. First, we do not assume access to a separate source-domain dataset; instead, we operate within a single large dataset that contains multiple related domains of interest. These domains can be viewed as micro-domains–related subdomains (e.g., ”Computers” and ”Clothing”) within a larger macro-domain (e.g., ”e-commerce Products”). In contrast, existing work on cross-domain EM [53, 54, 51] primarily focuses on transferring knowledge across distinct macro-domains. Second, we assume access to a limited amount of in-domain labeled data and study how to augment it under a constrained sample budget. Rather than transferring representations or model parameters between datasets, our formulation emphasizes selecting samples across domains within the same dataset.
@@ 3.2 Entity Matching
Given a filtered candidate set $N$, the goal of the matching phase is to determine which candidates contain matching entities. This is formulated as a binary classification task: for each candidate $(r_{1},r_{2})_{i}\in N$, we assign a label $y_{i}\in\{0,1\}$, where $y_{i}=1$ if the $r_{1}$ and $r_{2}$ are matching entities, and 0 otherwise. We refer to a candidate together with its corresponding label as a sample.
We adopt PLMs such as BERT [12] (or RoBERTa [28] and DistilBERT [44]) to perform the classification. A PLM refers to a transformer-based model trained on a large corpora of text, which can be fine-tuned for downstream tasks such as EM. Following the DITTO framework [27], we serialize an entity $r$ with $p$ attributes as:
@@ 3.3 Problem Definition
Consider the task of matching entities across two product domains: ”Shoes” and ”Jewelry.” For shoes, EM often depends on a small set of attributes with highly standardized values. For instance, two Nike Air Zoom Pegasus 40 running shoes (US size 10, Style Code DV3853-401, retail price \$130) can be matched with high confidence if their Brand, Size, and Style Code attributes align across both records. [^3] Please find the full item description [3]
@@ 4.1.1 The General Model (GEN)
The General Model (GEN) generates $X_{i}$ by randomly sampling from the candidate set, without regard to domain. For example, if we are training a model to match WDC ”Computers” pairs, we sample across all available domains (e.g., ”Electronics”, ”Clothing”, etc.) to fill the budget, including candidates from the target domain itself (see Example  1.1). GEN corresponds to the idea of training a single matcher across all domains, as discussed in Section 3.3. By randomly sampling across all domains, GEN simulates a unified model trained without regard to domain boundaries, while still respecting the budget constraint. GEN is motivated by the hypothesis that broad domain coverage may expose the model to a more diverse range of attribute patterns, potentially improving its generalization to unseen data. In practice, GEN represents the most domain-agnostic baseline.
@@ 4.3.1 The Ensemble of Models
Once several models are trained, their predictions can be aggregated to form a more accurate consensus. Given $k$ trained models, let $q\in[1,k]$ index a model. We assume access to a validation set $v_{i}$ for each domain $i\in\mathcal{D}$ and measure the validation performance of each model as $F1_{\text{VAL},q}$. We adopt a weighted soft voting scheme, in which the confidence outputs $Conf_{q}$ from each model are combined using weights proportional to their validation F1 scores. This provides a principled way to emphasize models that perform best during validation. Formally, the final ensemble confidence $Conf_{f}$ for a given candidate pair is computed as:

### 17 Mar 2026 v2

@@ Abstract
Entity Matching (EM)–the task of determining whether two data records refer to the same real-world entity–is a core task in data integration. Recent advances in deep learning have set a new standard for EM, particularly through fine-tuning Pretrained Language Models (PLMs) and, more recently, Large Language Models (LLMs). However, fine-tuning typically requires large amounts of labeled data, which are expensive and time-consuming to obtain. In the context of e-commerce matching, labeling scarcity varies widely across domains, raising the question of how to intelligently train accurate domain-specific EM models with limited labeled data. In this work we assume users have only a limited amount of labels for a specific target domain but have access to labeled data from other domains. We introduce BEACON, a distribution-aware, budget-aware framework for low-resource EM across domains. BEACON leverages the insight that embedding representations of pairwise candidate matches can guide the effective selection of out-of-domain samples under limited in-domain supervision. We conduct extensive experiments across multiple domain-partitioned datasets derived from established EM benchmarks, demonstrating that BEACON consistently outperforms state-of-the-art methods under different training budgets.
@@ 1 Introduction
Figure 1 shows three entity pairs produced by blocking, with each pair comparing two product records from e-commerce datasets. The samples span multiple product categories and illustrate the types of candidate pairs that serve as inputs to EM models. A solid green line between records indicates a match, while a red dotted line indicates a non-match. Samples (a) and (c) represent matches, where both entities in the pair refer to the same product (though this is not known during matching). For example, sample (a) compares the same RAM stick, and sample (c) compares the same power supply. Conversely, sample (b) is a non-match: although both products are baseball caps from the same brand, they are distinct items. The record data is incomplete–as can be seen with missing “brand” values–highlighting the challenges of real-world data systems and EM in practice.
Figure 2: Entity matching samples from Figure 1 grouped by the product category domain. The goal is to select a pairwise sample to complement the selected training set for EM in the “Computers” product category.
@@
Suppose we aim to train a model for the “Computers” category. Figure 2 shows the same three pairwise samples from Figure 1, now labeled by product category. Note that any pairs spanning different product categories are automatically removed during blocking, allowing the remaining data to be partitioned into distinct category-specific groups. To build a training set for “Computers,” we first include all in-domain pairs, as they are most likely to be relevant to our model. In our toy example, only pair (a) belongs to the domain and is selected first. If room remains in the budget after including all in-domain samples, we need to decide which out-of-domain examples are most “beneficial.”
@@
  • Experimental Evaluation. We adapt EM PLMs (e.g., DITTO [27]), recent EM domain adaptation frameworks (e.g., MFSN [51]), recent low-resource EM methods (e.g., Battleship [19] and PromptEM
     [56]), and recent LLM approaches (e.g., Jellyfish [61, 62]) to establish strong baselines for EMAD, over which we conduct extensive experiments showing that BEACON outperforms approaches across different budget constraints. For this evaluation, we derive nine domain-partitioned datasets from WDC [38], as well as two additional datasets derived from the WDC Product Corpus [41] and Abt-Buy [25], which we also release to support future research on EMAD. We further provide Open-source Code and Data. [^2] https://github.com/nbpulsone/BEACON
@@ 2 Related Work
While our work shares the high-level goal of improving EM under domain shifts, there are two important distinctions that highlight the novelty of our problem setting. First, we do not assume access to a separate source-domain dataset; instead, we operate within a single large dataset that contains multiple related domains of interest. These domains can be viewed as micro-domains–related subdomains (e.g., “Computers” and “Clothing”) within a larger macro-domain (e.g., “e-commerce Products”). In contrast, existing work on cross-domain EM [53, 54, 51] primarily focuses on transferring knowledge across distinct macro-domains. Second, we assume access to a limited amount of in-domain labeled data and study how to augment it under a constrained sample budget. Rather than transferring representations or model parameters between datasets, our formulation emphasizes selecting samples across domains within the same dataset.
@@ 3.2 Entity Matching
Given a filtered candidate set $N$, the goal of the matching phase is to determine which candidates contain matching entities. This is formulated as a binary classification task: for each candidate $(r_{1},r_{2})_{i}\in N$, we assign a label $y_{i}\in\{0,1\}$, where $y_{i}=1$ if $r_{1}$ and $r_{2}$ are matching entities, and 0 otherwise. We refer to a candidate together with its corresponding label as a sample.
We adopt PLMs such as BERT [12] (or RoBERTa [28] and DistilBERT [44]) to perform the classification. A PLM refers to a transformer-based model trained on a large corpus of text, which can be fine-tuned for downstream tasks such as EM. Following the DITTO framework [27], we serialize an entity $r$ with $p$ attributes as:
@@ 3.3 Problem Definition
Consider the task of matching entities across two product domains: “Shoes” and “Jewelry.” For shoes, EM often depends on a small set of attributes with highly standardized values. For instance, two Nike Air Zoom Pegasus 40 running shoes (US size 10, Style Code DV3853-401, retail price \$130) can be matched with high confidence if their Brand, Size, and Style Code attributes align across both records. [^3] Please find the full item description [3]
@@ 4.1.1 The General Model (GEN)
The General Model (GEN) generates $X_{i}$ by randomly sampling from the candidate set, without regard to domain. For example, if we are training a model to match WDC “Computers” pairs, we sample across all available domains (e.g., “Electronics,” “Clothing,” etc.) to fill the budget, including candidates from the target domain itself (see Example  1.1). GEN corresponds to the idea of training a single matcher across all domains, as discussed in Section 3.3. By randomly sampling across all domains, GEN simulates a unified model trained without regard to domain boundaries, while still respecting the budget constraint. GEN is motivated by the hypothesis that broad domain coverage may expose the model to a more diverse range of attribute patterns, potentially improving its generalization to unseen data. In practice, GEN represents the most domain-agnostic baseline.
@@ 4.3.1 The Ensemble of Models
Once several models are trained, their predictions can be aggregated to form a more accurate consensus. Given $k$ trained models, let $q\in[1,k]$ index a model. We assume access to a validation set $v_{i}$ for each domain $i\in\mathcal{D}$ and measure the validation F1 performance of each model as $F1_{\text{VAL},q}$. We adopt a weighted soft voting scheme, in which the confidence outputs $Conf_{q}$ from each model are combined using weights proportional to their validation F1 scores. This provides a principled way to emphasize models that perform best during validation. Formally, the final ensemble confidence $Conf_{f}$ for a given candidate pair is computed as: